OpenStudy (anonymous):

solve by addition method 3x^2-4y^2=-13, 2x^2+5y^2=22

OpenStudy (amistre64):

3x^2 -4y^2 =-13 ;*2 2x^2 +5y^2 =22 ;*-3 6x^2 -8y^2 =-26 -6x^2 -15y^2 =-66 ------------------- 0 -23y^2 = -92 y = sqrt(92/23)

OpenStudy (amistre64):

y = +- sqrt(92/23)..sooo 3x^2 - 4(92)/23 = -13

OpenStudy (amistre64):

x^2 = -13 +4(9)/23 ------------ 3 x = -13 +4(9)/23 sqrt( ------------ ) 3

OpenStudy (amistre64):

sounds a bit off....maybe

OpenStudy (amistre64):

you got an answer key to go by?

OpenStudy (amistre64):

where do a hyperbola and a circle meet.....

OpenStudy (anonymous):

no i dont have an answer key